OpenStudy (anonymous):

6y(2y+1)=12y+8y(squared)+9

OpenStudy (aum):

6y(2y+1) = 12y + 8y^2 + 9 distribute on the left hand side: 12y^2 + 6y = 12y + 8y^2 + 9 subtract 8y^2: 4y^2 + 6y = 12y + 9 subtract 12y: 4y^2 - 6y = 9 subtract 9 4y^2 - 6y - 9 = 0 solve the quadratic equation.

OpenStudy (anonymous):

Thankyouu

OpenStudy (aum):

You are welcome.